OpenStudy (anonymous):

A car travels at an average speed of 68 miles per hour. How many miles does it travel in 3 hours and 30 minutes?

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

ganeshie8 (ganeshie8):

3 hours and 30 minutes = 3.5 hours

OpenStudy (anonymous):

If you help please show all work, im doing this without a calculator and i missed class yesterday due to doc appt. thanks!

ganeshie8 (ganeshie8):

use this formula for distance travelled :- distance travelled = speed x time

ganeshie8 (ganeshie8):

oh you had a doc appt... wat happened... hope you're doing fine nw....

OpenStudy (anonymous):

so id set it up like N=68 * 3.5 ? And my daughter has border line ammonia and she needed to come for check up and had to get 2 shots, shes doing better i dont have class till monday but everything is due sunday night so im trying to get it done while shes asleep lol

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

distance = speed x time remember this for rest of the life lol... this formula wil is so useful

OpenStudy (anonymous):

so n = 238.0

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

yay i got it right! Btw have u heard a website called Aleks.com? bcuz thats what i have to do and i have to learn to do it without a cal bcuz i cant use one for my final!

ganeshie8 (ganeshie8):

never heard of that before... is that like flvs ?

OpenStudy (uri):

68 x3.5 =238 miles

OpenStudy (uri):

Easy :P

ganeshie8 (ganeshie8):

ha its very easy :)

OpenStudy (anonymous):

no its a website to help u learn the material instead of using the book. and its basic college math its hard for me because if always had a calcualor lol

ganeshie8 (ganeshie8):

ohhk... calculator must be allowed from college level

OpenStudy (anonymous):

im sure i can get this but do i set it up? A light bulb consumes 18900 watt-hours in 5 days and 6 hours. How many watt-hours does it consume per day? would i set that up like n= 18900w/h *120h

OpenStudy (anonymous):

Nope not in my class, no cals aloud which really sucks i had to relearn my multiplcation talble!!!

ganeshie8 (ganeshie8):

lol go back to good old times of timetables... ;)

ganeshie8 (ganeshie8):

18900 watt-hours in 5 days and 6 hours.

OpenStudy (anonymous):

5d=120h +6h = 126h

ganeshie8 (ganeshie8):

lets convert everything to hours :- 18900 watt-hours in 5x24 hours + 6 hours. 18900 watt-hours in 126 hours.

ganeshie8 (ganeshie8):

you're brilliant !

ganeshie8 (ganeshie8):

we set it up like this :- \(\frac{18900}{126} = \frac{x}{24}\)

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

where did the 24 come from

ganeshie8 (ganeshie8):

good q :) read the q once

ganeshie8 (ganeshie8):

it asks us to find watts consumed per \(day\)

ganeshie8 (ganeshie8):

\(day\) = 24 hours assuming the bulb is in a house built on earth ;p

OpenStudy (anonymous):

ahh i see ok so heres where i am at